# Global weak solutions for the inverse mean curvature flow in the Heisenberg group

### v2

@@ 3. Construction of barriers
Let $1<p\leq 2$, $\varepsilon\in[0,1]$, and let $\Omega\subset\mathbb{H}^{n}$ be an open set with $C^{2}$-smooth boundary and bounded complement satisfying assumption assuption ($\mathbf{HP_{\Omega}}$) with parameter $R_{0}$ such that $B_{R_{0}}(g_{0})\subset\Omega^{c}$ and $\mathbb{H}^{n}\setminus\Omega\subset B_{\bar{R}}(g_{0})$ for some $R_{0}<2<\bar{R}$. There exist $C>0$ depending only on $R_{0}$ and $Q$ and there exists $\widehat{C}>1$ depending only on $R_{0}$, $\bar{R}$ and $Q$ such that for $v=v^{\varepsilon}_{p}$ the minimizer of (1.10) in the set $\dot{W}^{1,p}_{1,\varepsilon}(\Omega)$ we have
(5.1) $v(g)\geq\widehat{C}^{-\frac{\varepsilon^{4}}{p-1}}\left(\frac{\|g_{0}^{-1}*g\|}{R_{0}}\right)^{-\gamma}\qquad\hbox{on }\quad\overline{\Omega}\,,\qquad\hbox{with}\quad\gamma=\frac{Q-p}{p-1}\,,\quad\widehat{C}=\left(4\frac{\bar{R}}{R_{0}}\right)^{C}\,.$
@@
First notice that the claim is true for $\varepsilon=0$ by chosing any $\widehat{C}>0$. Indeed in that case both sides in (5.1) are continuous and finite energy $p$-harmonic functions in $\Omega$ with respect to the sub-Riemannian metric $|\cdot|_{0}$, moreover they are clearly well ordered on $\partial\Omega$, so that inequality (5.1) holds in the whole $\Omega$ as it follows from the comparison principle for finite energy $p$-harmonic functions.
Assuming $\varepsilon\in(0,1]$ in the sequel, up to traslations we may also assume $g_{0}=0$. Notice that the right hand side in (5.1) is no longer $p$-harmonic in $\Omega$ but a slight tilt in the exponent still gives a subsolution according to Proposition 3.3. Then we iterate this property at different scales and with different tilts in the exponent in order to prove that (5.1) actually holds for suitably chosen $\widehat{C}>0$.
In what follows for $\varepsilon\in(0,1]$ we set $\mu(\varepsilon)=\frac{\varepsilon^{4}}{p-1}\left(\frac{8}{R_{0}^{2}}+\frac{32Q}{R_{0}^{4}}\right)$, so that in view of Lemma 3.2 and Proposition 3.3 the function $\|\cdot\|^{-(\gamma+\mu(\varepsilon))}$ has finite energy and it satisfies (3.33), i.e., it is weakly $p$-subharmonic w.r.to the metric $|\cdot|_{\varepsilon}$ outside the ball $B_{R_{0}}(0)$, whence comparison principle yields $R_{0}^{-\gamma-\mu(\varepsilon)}v\geq\|\cdot\|^{-(\gamma+\mu(\varepsilon))}$ in $\overline{\Omega}$. Now for $\bar{R}>2>R_{0}$ such that $\mathbb{H}^{n}\setminus\Omega\subset B_{\bar{R}}(0)$, if $\widetilde{\Omega}=\mathbb{H}^{n}\setminus\overline{B_{\bar{R}}(0)}$ then we have $v\in\dot{W}^{1,p}_{\varepsilon}(\widetilde{\Omega})\cap C(\overline{\widetilde{\Omega}})$. Next, using homogeneous dilations we define by induction
@@
Clearly $\mu(\varepsilon_{j})\leq\frac{C\varepsilon^{4}}{p-1}2^{-j-1}$ for all $j\geq 0$, where $C>0$ depends only on $R_{0}$ and $Q$, whence $\sum_{j}\mu(\varepsilon_{j})\leq\frac{C\varepsilon^{4}}{p-1}$, $m_{j}\geq m_{0}\left(2^{C}\right)^{\varepsilon^{4}/(1-p)}$ and $v_{j}\geq m_{j}\|\cdot\|^{-(\gamma+\mu(\varepsilon_{j}))}\geq m_{0}\left(4^{C}\right)^{\varepsilon^{4}/(1-p)}\|\cdot\|^{-\gamma}$ on $B_{1}(0)^{c}\cap\overline{B_{2}(0)}$.
@@
$m_{0}=\bar{R}^{\gamma}\,\min_{\partial B_{\bar{R}}(0)}\,v\geq\bar{R}^{\gamma}\left(\frac{\bar{R}}{R_{0}}\right)^{-\gamma-\mu(\varepsilon)}\geq R_{0}^{\gamma}\left(\frac{\bar{R}}{R_{0}}\right)^{-\frac{C\varepsilon^{4}}{p-1}}$
@@
$v_{j}(g)\geq R_{0}^{\gamma}\left(\frac{\bar{R}}{R_{0}}\right)^{-\frac{C\varepsilon^{4}}{p-1}}\left(4^{C}\right)^{\varepsilon^{4}/(1-p)}\|g\|^{-\gamma}=\left(4\frac{\bar{R}}{R_{0}}\right)^{-\frac{C\varepsilon^{4}}{p-1}}\left(\frac{\|g\|}{R_{0}}\right)^{-\gamma}\quad\hbox{for each}\quad g\in B_{1}(0)^{c}\cap\overline{B_{2}(0)}\,,$
therefore the identity $v_{j}(g)=\bar{R}^{\gamma}_{j}v(\delta_{\bar{R}_{j}}(g))$ for any $j\in\mathbb{N}$ yield inequality (5.1) on $\mathbb{H}^{n}\setminus B_{\bar{R}}(0)$ by chosing $\widehat{C}=\left(4\frac{\bar{R}}{R_{0}}\right)^{C}>1$. Finally, since $R_{0}^{-\gamma-\mu(\varepsilon)}v\geq\|\cdot\|^{-(\gamma+\mu(\varepsilon))}$ in $\overline{\Omega}$ yields $R_{0}^{-\gamma}v\geq(4\bar{R}/R_{0})^{-\mu(\varepsilon)}\|\cdot\|^{-\gamma}$ in $\overline{\Omega}\cap B_{\bar{R}}(0)$, then from the choice of $C$ above and the corresponding definition of $\widehat{C}$ inequality (5.1) actually holds in the whole $\overline{\Omega}$. This closes the proof. ∎
In order to obtain a pointwise upper bound for the $p$-capacitary potentials we start with the following auxiliary result announced in Section 2 about the Riemannian perimeters $Per_{\varepsilon}(\cdot)$ and their relation with the horizontal perimeter $Per_{0}(\cdot)$. The reader should refer to [33, Section 5] for a similar statement, although for a slightly different definition of $Per_{\varepsilon}(\cdot)$ corresponding to a family of vector fields different from $\{X_{i},Y_{i},T_{\varepsilon}\}$.
@@
Let $\widetilde{\Omega}\subset\mathbb{H}^{n}$ be a bounded open set with $C^{1}$-smooth boundary, $\mathcal{H}^{2n}$ be the Eucledean $2n$-dimensional Hausdorff measure and let $\hat{n}$ be the Euclidean outer normal vector field.
@@
As a consequence $\varepsilon\to\varepsilon Per_{\varepsilon}(\widetilde{\Omega})$ is non decreasing, $\varepsilon Per_{\varepsilon}(\widetilde{\Omega})\leq Per_{1}(\widetilde{\Omega})$ for any $\varepsilon\in(0,1]$, and $\varepsilon Per_{\varepsilon}(\widetilde{\Omega})\to Per_{0}(\widetilde{\Omega})$ as $\varepsilon\to 0$.
@@
It is clearly enough to establish the validity of (5.3), whence the the other properties follow easily taking into account the representation formula for $P_{0}(\widetilde{\Omega})$ already proved in [12, Section 3] (see also [33] and [21, Section 7]).
The proof of (5.3) is very similar to the one in [12, Section 3] for $Per_{0}(\widetilde{\Omega})$. Indeed, recall that for any smooth vector field $Z_{\varphi}=\sum_{i=1}^{n}\left[\varphi_{i}X_{i}+\varphi_{n+i}Y_{i}\right]+\varphi_{2n+1}T_{\varepsilon}$ associated to a smooth map $\varphi:\mathbb{H}^{n}\to\mathbb{R}^{2n+1}$ its Riemannian divergence is given by ${\rm div}_{\varepsilon}Z_{\varphi}=\sum_{i=1}^{n}\left[(X_{i}\varphi_{i}+Y_{i}\varphi_{n+i}\right]+T_{\varepsilon}\varphi_{2n+1}\,,$ and moreover that we have the identities $X_{i}\varphi_{i}=\partial_{x_{i}}\varphi_{i}-\partial_{t}(\frac{y_{i}}{2}\varphi_{i})$, $Y_{i}\varphi_{i}=\partial_{y_{i}}\varphi_{n+i}+\partial_{t}(\frac{x_{i}}{2}\varphi_{n+i})$ and $T_{\varepsilon}\varphi_{2n+1}=\partial_{t}(\varepsilon\varphi_{2n+1})$.
@@
[…]
@@
As a consequence, if $\Omega$ satisfies also assumption $(\bf{HP_{\Omega}})$ with parameter $R_{0}>0$ then there exist $C^{\prime}>0$ depending only on $Q$ and $R_{0}$ such that $\|v\|_{L^{\sigma,\infty}(\Omega)}\leq C^{\prime}$.
@@
On the other hand, since $\widetilde{\Omega}$ has smooth boundary $\partial\widetilde{\Omega}=\partial\Omega$, the same holds for $\widetilde{\Omega}_{r,\varepsilon}=\{v>1-r\}$ for any $0<r<\tau$ and for $\tau$ small enough because $\nabla_{\varepsilon}v\neq 0$ on $\partial\Omega$. As a consequence $Per_{\varepsilon}(\widetilde{\Omega}_{r,\varepsilon})=(1+o(1))Per_{\varepsilon}(\widetilde{\Omega})$ as $r\to 0$. Since $\mathcal{L}^{n+1}=\varepsilon\,dvol_{\varepsilon}(\cdot)$, then the Riemannian coarea formula together with the previous expansions yield
@@
(5.6) $\int|\nabla_{\varepsilon}\min\{v,\rho\}|_{\varepsilon}^{p}=\int_{\{v<\rho\}}|\nabla_{\varepsilon}v|_{\varepsilon}^{p}\leq\rho\||\nabla_{\varepsilon}v|_{\varepsilon}\|_{L^{\infty}(\partial\Omega)}^{p-1}Per_{1}(\widetilde{\Omega})\,,$
@@
$\rho^{p}|\{v>\rho\}|^{p/p^{*}}\leq\left(\int(\min\{v,\rho\})^{p^{*}}\right)^{p/p^{*}}\leq(Cp^{*})^{p}\rho\||\nabla_{\varepsilon}v|_{\varepsilon}\|_{L^{\infty}(\partial\Omega)}^{p-1}Per_{1}(\widetilde{\Omega})\,,$
@@
and the conclusion follows as the right hand side is uniformly bounded for $p\in(1,2]$. ∎
@@
Let $1<p\leq 2$, $p_{*}=\frac{Q(p-1)}{Q-p/2}$ and for $r>\varepsilon>0$ let us set $\Omega_{r}=B^{\varepsilon}_{5r}(\bar{g})\setminus\overline{B^{\varepsilon}_{r}(\bar{g})}$ and $\widetilde{\Omega}_{r}=B^{\varepsilon}_{4r}(\bar{g})\setminus\overline{B^{\varepsilon}_{2r}(\bar{g})}\subset\Omega_{r}$ There exists $C_{*}>1$ depending only on $Q$ such that if $v=v^{\varepsilon}_{p}\in W^{1,p}_{\varepsilon}(\Omega_{r})$ is nonegative and $p$-harmonic in $\Omega_{r}$ w.r. to the metric $|\cdot|^{2}_{\varepsilon}$ then for any $q\geq p_{*}$ we have
@@
As in the proof of Proposition 4.2 we first derive a Caccioppoli-type inequality for suitable powers of $v$. Thus, for $\beta>-1/p$ and for nonnegative $d_{\varepsilon}$-Lipschitz function $\zeta$ compactly supported in $\Omega_{r}$ to be specified later, we test the equation for $v$ with $\zeta^{p}v^{1+\beta p}$, hence integration by parts, Hölder and Young inequalities combined as in the proof of (4.18) yield
@@
Next, in analogy with the previous proof we specialize inequality (5.8) by chosing suitable test functions defined in terms of the metric $d_{\varepsilon}$ in order to obtain (5.7) from (5.8) through Moser iteration. For $\frac{r}{2}<\frac{1}{2}\rho_{2}\leq\rho_{1}<\rho_{2}\leq 2r$ we consider $h:[0,\infty)\to[0,1]$ a Lipschitz and piecewhise linear function, with $\{h\equiv 1\}=[3r-\rho_{1},3r+\rho_{1}]$, $\{h\equiv 0\}=[0,3r-\frac{\rho_{1}+\rho_{2}}{2}]\cup[3r+\frac{\rho_{1}+\rho_{2}}{2},\infty)$, and $h$ is linear otherwise. Hence, in view of the inequality $|\nabla_{\varepsilon}d_{\varepsilon}(\cdot,\bar{g})|_{\varepsilon}\leq 1$ on $\mathbb{H}^{n}$ the function $\zeta(g)=h(d_{\varepsilon}(g,0))$ by construction is $d_{\varepsilon}$-Lipschitz on $\mathbb{H}^{n}$ with $\zeta\equiv 1$ on $\overline{B^{\varepsilon}_{3r+\rho_{1}}}\setminus B^{\varepsilon}_{3r-\rho_{1}}$, $spt\,\zeta\subset\subset B^{\varepsilon}_{3r+\rho_{2}}\setminus\overline{B^{\varepsilon}_{3r-\rho_{2}}}$, and $|\nabla_{\varepsilon}\zeta|_{\varepsilon}\leq\frac{2}{\rho_{2}-\rho_{1}}\,$ a.e. in $\mathbb{H}^{n}$.
@@
Next, using (5.9) we perform Moser iteration with exponents $\{\beta_{i}\}_{i\geq 0}$ where $\beta_{0}>-1/p$ is chosen such that $(1+\beta_{0})p=p_{*}$ and $\beta_{i+1}=p/(Q-p)+\beta_{i}Q/(Q-p)$, so that $\{\beta_{i}\}\subset(-1/p,\infty)$ and it is increasing to infinity. Indeed, note that if $s=p(\beta+1)$ then $\beta>-1/p$ iff $s>p-1$ so that for $s_{i}=p(\beta_{i}+1)$ we have $s_{i+1}=s_{i}Q/(Q-p)=p_{*}\frac{Q^{i+1}}{(Q-p)^{i+1}}$ is increasing and $s_{i}>(p-1)\frac{Q^{i+1}}{(Q-1/2)^{i+1}}\uparrow+\infty$ as $i\to\infty$. Moreover, since $\beta\to\frac{(1+\beta)p}{1+\beta p}$ is decreasing for $\beta>-1/p$ one has $\frac{(1+\beta_{i})p}{1+\beta_{i}p}\leq\frac{(1+\beta_{0})p}{1+\beta_{0}p}\leq 2Q$ for any $i\geq 0$ and $p>1$.
@@
Let $1<p\leq 2$, $\varepsilon\in[0,1]$, and let $\Omega\subset\mathbb{H}^{n}$ be an open set with $C^{2}$-smooth boundary and bounded complement satisfying assumption assuption ($\mathbf{HP_{\Omega}}$) with parameter $R_{0}$ such that $B_{R_{0}}(g_{0})\subset\Omega^{c}$ and $\mathbb{H}^{n}\setminus\Omega\subset B_{\bar{R}}(g_{0})$ for some $R_{0}<2<\bar{R}$. There exist $C_{0}>1$ depending only on $R_{0}$, $\bar{R}$, $\partial\Omega$ and $Q$ such that for $v=v^{\varepsilon}_{p}$ the minimizer of (1.10) in the set $\dot{W}^{1,p}_{1,\varepsilon}(\Omega)$ we have
@@
On the other hand, notice that still Lemma 2.2 gives $2C\varepsilon<C2^{i}\bar{R}\leq d_{\varepsilon}(g_{0},g)\leq C\|g_{0}^{-1}*g\|_{\varepsilon}$, whence $\|g_{0}^{-1}*g\|\leq 2\|g_{0}^{-1}*g\|_{\varepsilon}\leq 2Cd_{\varepsilon}(g_{0},g)\leq 8C^{2}2^{i}\bar{R}=R_{0}\tilde{r}_{i}$ and in turn $v(g)\geq\widehat{C}^{1/(1-p)}\tilde{r}_{i}^{-\gamma}$ on $\Omega_{r_{i}}$, where $\tilde{r}_{i}=8C^{2}2^{i}\bar{R}/R_{0}=8Cr_{i}/R_{0}$, because of Proposition 5.1.

### v3

@@ 3. Construction of barriers
Let $1<p\leq 2$, $\varepsilon\in[0,1]$, and let $\Omega\subset\mathbb{H}^{n}$ be an open set with $C^{2}$-smooth boundary and bounded complement satisfying assumption assumption ($\mathbf{HP_{\Omega}}$) with parameter $R_{0}$ such that $B_{R_{\ast}}(g_{0})\subset\Omega^{c}$ and $\mathbb{H}^{n}\setminus\Omega\subset B_{\bar{R}}(g_{0})$ for some $R_{\ast}<2<\bar{R}$. There exist $C>0$ depending only on $R_{\ast}$ and $Q$ and there exists $\widehat{C}>1$ depending only on $_{\ast}, R¯\bar{R}$ and $Q$ such that for $v=v^{\varepsilon}_{p}$ the minimizer of (1.10) in the set $\dot{W}^{1,p}_{1,\varepsilon}(\Omega)$ we have
(5.1) $v(g)\geq\widehat{C}^{-\frac{\varepsilon^{4}}{p-1}}\left(\frac{\|g_{0}^{-1}*g\|}{R_{\ast}}\right)^{-\gamma}\qquad\hbox{on }\quad\overline{\Omega}\,,\qquad\hbox{with}\quad\gamma=\frac{Q-p}{p-1}\,,\quad\widehat{C}=\left(4\frac{\bar{R}}{R_{\ast}}\right)^{C}\,.$
@@
First notice that the claim is true for $\varepsilon=0$ by choosing any $\widehat{C}>0$. Indeed in that case both sides in (5.1) are continuous and finite energy $p$-harmonic functions in $\Omega$ with respect to the sub-Riemannian metric $|\cdot|_{0}$, moreover they are clearly well ordered on $\partial\Omega$, so that inequality (5.1) holds in the whole $\Omega$ as it follows from the comparison principle for finite energy $p$-harmonic functions.
Assuming $\varepsilon\in(0,1]$ in the sequel, up to translations we may also assume $g_{0}=0$. Notice that the right hand side in (5.1) is no longer $p$-harmonic in $\Omega$ but a slight tilt in the exponent still gives a subsolution according to Proposition 3.3. Then we iterate this property at different scales and with different tilts in the exponent in order to prove that (5.1) actually holds for suitably chosen $\widehat{C}>0$.
[…]
@@
Clearly $\mu(\varepsilon_{j})\leq\frac{C\varepsilon^{4}}{p-1}2^{-j-1}$ for all $j\geq 0$, where $C>0$ depends only on $R_{\ast}$ and $Q$, whence $\sum_{j}\mu(\varepsilon_{j})\leq\frac{C\varepsilon^{4}}{p-1}$, $m_{j}\geq m_{0}\left(2^{C}\right)^{\varepsilon^{4}/(1-p)}$ and $v_{j}\geq m_{j}\|\cdot\|^{-(\gamma+\mu(\varepsilon_{j}))}\geq m_{0}\left(4^{C}\right)^{\varepsilon^{4}/(1-p)}\|\cdot\|^{-\gamma}$ on $B_{1}(0)^{c}\cap\overline{B_{2}(0)}$.
@@
$m_{0}=\bar{R}^{\gamma}\,\min_{\partial B_{\bar{R}}(0)}\,v\geq\bar{R}^{\gamma}\left(\frac{\bar{R}}{R_{\ast}}\right)^{-\gamma-\mu(\varepsilon)}\geq R_{\ast}^{\gamma}\left(\frac{\bar{R}}{R_{\ast}}\right)^{-\frac{C\varepsilon^{4}}{p-1}}$
@@
$v_{j}(g)\geq R_{\ast}^{\gamma}\left(\frac{\bar{R}}{R_{\ast}}\right)^{-\frac{C\varepsilon^{4}}{p-1}}\left(4^{C}\right)^{\varepsilon^{4}/(1-p)}\|g\|^{-\gamma}=\left(4\frac{\bar{R}}{R_{\ast}}\right)^{-\frac{C\varepsilon^{4}}{p-1}}\left(\frac{\|g\|}{R_{\ast}}\right)^{-\gamma}\quad\hbox{for each}\quad g\in B_{1}(0)^{c}\cap\overline{B_{2}(0)}\,$,
therefore the identity $v_{j}(g)=\bar{R}^{\gamma}_{j}v(\delta_{\bar{R}_{j}}(g))$ for any $j\in\mathbb{N}$ yield inequality (5.1) on $\mathbb{H}^{n}\setminus B_{\bar{R}}(0)$ by choosing $\widehat{C}=\left(4\frac{\bar{R}}{R_{\ast}}\right)^{C}>1$. Finally, since $R_{\ast}^{-\gamma-\mu(\varepsilon)}v\geq\|\cdot\|^{-(\gamma+\mu(\varepsilon))}$ in $\overline{\Omega}$ yields $R_{\ast}^{-\gamma}v\geq(4\bar{R}/R_{\ast})^{-\mu(\varepsilon)}\|\cdot\|^{-\gamma}$ in $\overline{\Omega}\cap B_{\bar{R}}(0)$, then from the choice of $C$ above and the corresponding definition of $\widehat{C}$ inequality (5.1) actually holds in the whole $\overline{\Omega}$. This closes the proof. ∎
In order to obtain a pointwise upper bound for the $p$-capacitary potentials we start with the following auxiliary result announced in Section 2 about the Riemannian perimeters $Per_{\varepsilon}(\cdot)$ and their relation with the horizontal perimeter $Per_{0}(\cdot)$. The reader should refer to [35, Section 5] for a similar statement, although for a slightly different definition of $Per_{\varepsilon}(\cdot)$ corresponding to a family of vector fields different from $\{X_{i},Y_{i},T_{\varepsilon}\}$.
@@
Let $\widetilde{\Omega}\subset\mathbb{H}^{n}$ be a bounded open set with $C^{1}$-smooth boundary, $\mathcal{H}^{2n}$ be the Euclidean $2n$-dimensional Hausdorff measure and let $\hat{n}$ be the Euclidean outer normal vector field.
@@
(5.3) $\varepsilon Per_{\varepsilon}(\widetilde{\Omega})=\int_{\partial\widetilde{\Omega}}\left(\sum_{i=1}^{n}\left[\left(X_{i}\cdot\hat{n}\right)^{2}+\left(Y_{i}\cdot\hat{n}\right)^{2}\right]+\varepsilon^{2}\left(T\cdot\hat{n}\right)^{2}\right)^{1/2}\,d\mathcal{H}^{2n}\,.$
@@
[…]
@@
$\varepsilon\int_{\widetilde{\Omega}}{\rm div}_{\varepsilon}Z_{\varphi}\,dvol_{\varepsilon}=\int_{\partial\widetilde{\Omega}}\left(\sum_{i=1}^{n}\left[\varphi_{i}X_{i}\cdot\hat{n}+\varphi_{n+i}Y_{i}\cdot\hat{n}\right]+\varphi_{2n+1}\varepsilon T\cdot\hat{n}\right)\,d\mathcal{H}^{2n}\,$,
@@
As a consequence, if $\Omega$ satisfies also assumption $(\bf{HP_{\Omega}})$ with parameter $R_{0}>0$ then there exist $C^{\prime}>0$ depending only on $Q$, $R_{0}$ and $\Omega$ such that $\|v\|_{L^{\sigma,\infty}(\Omega)}\leq C^{\prime}$.
@@
$\int_{\{v<\rho\}}|\nabla_{\varepsilon}v|_{\varepsilon}^{p}$ $\leq(1+o(1))\||\nabla_{\varepsilon}v|_{\varepsilon}\|_{L^{\infty}(\partial\Omega)}^{p-1}\frac{\rho}{\tau}\int_{\{1-\tau<v\}}|\nabla_{\varepsilon}v|_{\varepsilon}\varepsilon dvol_{\varepsilon}$
@@
[…]
@@
[…]
@@
and the conclusion follows as the right hand side is uniformly bounded for $p\in(1,2]$ by $C^{\prime}>0$ with the claimed dependence of the parameters. ∎
@@
Let $1<p\leq 2$, $p_{*}=\frac{Q(p-1)}{Q-p/2}$ and for $r>\varepsilon>0$ let us set $\Omega_{r}=B^{\varepsilon}_{5r}(\bar{g})\setminus\overline{B^{\varepsilon}_{r}(\bar{g})}$ and $\widetilde{\Omega}_{r}=B^{\varepsilon}_{4r}(\bar{g})\setminus\overline{B^{\varepsilon}_{2r}(\bar{g})}\subset\Omega_{r}$ . There exists $C_{*}>1$ depending only on $Q$ such that if $v=v^{\varepsilon}_{p}\in W^{1,p}_{\varepsilon}(\Omega_{r})$ is nonegative and $p$-harmonic in $\Omega_{r}$ w.r. to the metric $|\cdot|^{2}_{\varepsilon}$ then for any $q\geq p_{*}$ we have
@@
As in the proof of Proposition 4.2 we first derive a Caccioppoli-type inequality for suitable powers of $v$. Thus, for $\beta>-1/p$ and for nonnegative $d_{\varepsilon}$-Lipschitz function $\zeta$ compactly supported in $\Omega_{r}$ to be specified later, we test the equation for $v$ with $\zeta^{p}v^{1+\beta p}$, hence integration by parts, Hölder and Young inequalities combined as in the proofs of (4.18) and (4.19) yield
@@
Next, in analogy with the previous proof we specialize inequality (5.8) by choosing suitable test functions defined in terms of the metric $d_{\varepsilon}$ in order to obtain (5.7) from (5.8) through Moser iteration. For $\frac{r}{2}<\frac{1}{2}\rho_{2}\leq\rho_{1}<\rho_{2}\leq 2r$ we consider $h:[0,\infty)\to[0,1]$ a Lipschitz and piecewise linear function, with $\{h\equiv 1\}=[3r-\rho_{1},3r+\rho_{1}]$, $\{h\equiv 0\}=[0,3r-\frac{\rho_{1}+\rho_{2}}{2}]\cup[3r+\frac{\rho_{1}+\rho_{2}}{2},\infty)$, and $h$ is linear otherwise. Hence, in view of the inequality $|\nabla_{\varepsilon}d_{\varepsilon}(\cdot,\bar{g})|_{\varepsilon}\leq 1$ on $\mathbb{H}^{n}$ the function $\zeta(g)=h(d_{\varepsilon}(g,0))$ by construction is $d_{\varepsilon}$-Lipschitz on $\mathbb{H}^{n}$ with $\zeta\equiv 1$ on $\overline{B^{\varepsilon}_{3r+\rho_{1}}}\setminus B^{\varepsilon}_{3r-\rho_{1}}$, $spt\,\zeta\subset\subset B^{\varepsilon}_{3r+\rho_{2}}\setminus\overline{B^{\varepsilon}_{3r-\rho_{2}}}$, and $|\nabla_{\varepsilon}\zeta|_{\varepsilon}\leq\frac{2}{\rho_{2}-\rho_{1}}\,$ a.e. in $\mathbb{H}^{n}$.
@@
Next, using (5.9) we perform Moser iteration with exponents $\{\beta_{i}\}_{i\geq 0}$ where $\beta_{0}>-1/p$ is chosen such that $(1+\beta_{0})p=p_{*}$ and $\beta_{i+1}=p/(Q-p)+\beta_{i}Q/(Q-p)$, so that $\{\beta_{i}\}\subset(-1/p,\infty)$ and it is increasing to infinity. Indeed, note that if $s=p(\beta+1)$ then $\beta>-1/p$ iff s>p1s>p-1 so that for $s_{i}=p(\beta_{i}+1)$ we have $s_{i+1}=s_{i}Q/(Q-p)=p_{*}\frac{Q^{i+1}}{(Q-p)^{i+1}}$ is increasing and $s_{i}>(p-1)\frac{Q^{i+1}}{(Q-1/2)^{i+1}}\uparrow+\infty$ as $i\to\infty$. Moreover, since $\beta\to\frac{(1+\beta)p}{1+\beta p}$ is decreasing for $\beta>-1/p$ one has $\frac{(1+\beta_{i})p}{1+\beta_{i}p}\leq\frac{(1+\beta_{0})p}{1+\beta_{0}p}\leq 2Q$ for any $i\geq 0$ and $p>1$.
@@
Let $1<p\leq 2$, $\varepsilon\in[0,1]$, and let $\Omega\subset\mathbb{H}^{n}$ be an open set with $C^{2}$-smooth boundary and bounded complement satisfying assumption assumption ($\mathbf{HP_{\Omega}}$) with parameter $R_{0}$ such that $B_{R_{\ast}}(g_{0})\subset\Omega^{c}$ and $\mathbb{H}^{n}\setminus\Omega\subset B_{\bar{R}}(g_{0})$ for some $R_{\ast}<2<\bar{R}$. There exist $C_{0}>1$ depending only on $Q$, $_{0}, RR_{\ast}$, $\bar{R}$ and $\partial\Omega$ such that for $v=v^{\varepsilon}_{p}$ the minimizer of (1.10) in the set $\dot{W}^{1,p}_{1,\varepsilon}(\Omega)$ we have
@@
On the other hand, notice that still Lemma 2.2 gives $2C\varepsilon<C2^{i}\bar{R}\leq d_{\varepsilon}(g_{0},g)\leq C\|g_{0}^{-1}*g\|_{\varepsilon}$, whence $\|g_{0}^{-1}*g\|\leq 2\|g_{0}^{-1}*g\|_{\varepsilon}\leq 2Cd_{\varepsilon}(g_{0},g)\leq 8C^{2}2^{i}\bar{R}=R_{\ast}\tilde{r}_{i}$ and in turn $v(g)\geq\widehat{C}^{1/(1-p)}\tilde{r}_{i}^{-\gamma}$ on $\Omega_{r_{i}}$, where $\tilde{r}_{i}=8C^{2}2^{i}\bar{R}/R_{\ast}=8Cr_{i}/R_{\ast}$, because of Proposition 5.1.